### v1

# Open Problems, and Future Trends

Abstract
As an emerging technology that enables seamless integration between the physical and virtual worlds, the Metaverse has great potential to be deployed in the industrial production field with the development of extended reality (XR) and next-generation communication networks. This deployment, called the Industrial Metaverse, is used for product design, production operations, industrial quality inspection, and product testing. However, there lacks of in-depth understanding of the enabling technologies associated with the Industrial Metaverse. This encompasses both the precise industrial scenarios targeted by each technology and the potential migration of technologies developed in other domains to the industrial sector. Driven by this issue, in this article, we conduct a comprehensive survey of the state-of-the-art literature on the Industrial Metaverse. Specifically, we first analyze the advantages of the Metaverse for industrial production. Then, we review a collection of key enabling technologies of the Industrial Metaverse, including blockchain (BC), digital twin (DT), 6G, XR, and artificial intelligence (AI), and analyze how these technologies can support different aspects of industrial production. Subsequently, we present numerous formidable challenges encountered within the Industrial Metaverse, including confidentiality and security concerns, resource limitations, and interoperability constraints. Furthermore, we investigate the extant solutions devised to address them. Finally, we briefly outline several open issues and future research directions of the Industrial Metaverse.
Index Terms:
Metaverse, industrial, enabling technologies, BC, DT, 6G, XR, AI.
I Introduction
The Metaverse is a nascent concept that presents a model of interconnection between the virtual and real worlds[1]. This digital system integrates both worlds into a cohesive system, providing users with a high degree of freedom to create content and modify modules as they wish, while also allowing different participants to define their own expressions. Some scholars define the Metaverse as a collection of virtual time and space, which leverages technologies such as virtual reality (VR) and augmented reality (AR). People continuously generate new content through smart wearables as inputs. It is also considered to be the latest phase of visual immersion technology, which has applications in all areas of production and life as an online digital space. Although there is no precise definition of the term “Metaverse” at this stage, its underlying technical system has matured and has been successfully applied in fields such as education, healthcare, industrial manufacturing, among others. The Industrial Metaverse refers to the establishment of a shared virtual space empowered by Metaverse technologies, to support multi-user, multi-device industrial scenarios for 3D modeling and immersive interaction.
Especially in the context of the COVID-19 pandemic, which has led to a widespread shift towards remote work, and the continuous advancement of digital technologies, many production activities that traditionally required physical presence can be transformed to enable employees to operate online, similar to a digital twin (DT). However, the difference is that DT is only as a digital replica of the physical elements of the real world, realizing the mapping from the physical world to the digital world. In contrast, Industrial Metaverse is a collective virtual shared space that contains all virtual worlds and derived elements, and virtual space digital individuals for online sharing, interactive perception and content generation. This feature is particularly suited to industrial scenarios, such as detecting part performance through nearly costless fault simulations, creating experimental conditions that cannot be replicated in reality, and expanding operational capabilities through the linkage between the virtual and physical worlds. This can guide physical industrial operations, reduce costs, and improve efficiency by empowering all aspects of the industrial process. Compared to typical virtual worlds, the emphasis of an industrial virtual world lies in simulating real-world industrial processes, manufacturing environments, or operational scenarios. The Industrial Metaverse places a higher demand on cost control for industrial manufacturing, requiring its design and implementation to consider cost efficiency and optimization, aiming to minimize deployment and maintenance costs. Additionally, the Industrial Metaverse involves collaboration among different enterprises in aspects such as industrial production and supply chain management, thus requiring a high level of trust in the execution of interactions. The key benefits of Metaverse are highlighted as follows:
  • Low-cost Simulation: The Metaverse can be used to simulate production processes in a cost-effective way. It can be used to implement various services by defining virtualized roles and scenarios, to create an orderly hierarchy in an enterprise virtual environment, and to generate content and simulation results that can be used directly to predict and optimize the situation in a real factory.
  • Cross-regional Collaboration: The Metaverse transcends geographical constraints and acquires data from diverse sensors and production lines, empowering manufacturers or disparate departments within the same manufacturing entity, geographically scattered, to collaborate on production, thereby enhancing production efficiency.
  • Secure Interaction Assurance: Through the integration of security technologies such as Non-Fungible Tokens (NFTs), the Industrial Metaverse can achieve secure protection of digital ownership, thereby aiding enterprises in preventing unauthorized user intrusions and data leaks during collaborative production processes.
    Figure 1: The development in the Metaverse from the perspective of technological updates. The concept of the Metaverse can be traced back to the VR concept proposed by Jaron Lanier in 1984. The formal concept of the Metaverse, however, was introduced in 1994. Over the following 30 years, enabling technologies and industries associated with the Metaverse (particularly the gaming industry) have continued to evolve. It was not until 2021 that there was a rapid surge in research achievements related to the Metaverse, which is why 2021 is also referred to as the “Metaverse Year”.
In summary, the Industrial Metaverse, as a novel intelligent manufacturing paradigm, is poised to catalyze an unprecedented industrial upgrading, and its inherent value is expected to significantly exceed its consumption value. Fig. 1 shows the development of the main enabling technologies in the Metaverse.
Sure, here’s the same table without the blue font color commands:
TABLE I: Key Abbreviations.
Abbreviation Explanation
BC Blockchain
DT Digital Twin
XR Extended Reality
IoT Internet of Things
IIoT Industrial Internet of Things
XRI XR-IoT
VR Virtual Reality
AR Augmented Reality
MR Mixed Reality
MSF Metaverse Standards Forum
SVM Support Vector Machines
LSTM Long Short-Term Memory
GANs Generative Adversarial Networks
PT Physical Twin
FL Federated Learning
VSPs Virtual Service Providers
AI Artificial Intelligence
ASP AI Service Provider
MEC Mobile Edge Computing
NFT Non-Fungible Token
EIoT Energy Internet of Things
HCI Human-computer Interaction
O&M Operations & Maintenance
BIM Building Information Modeling
PLC Product Lifecycle
CMS Cyber-Manufacturing Systems
HMDs Head-mounted Displays
DeAI Decentralized AI
CeAI Centralized AI
AIGC AI Generated Content
URLLC Ultra-reliable and Low-latency Communication
Here’s the table without the blue font color commands:
TABLE II: Survey papers of Industrial Metaverse.
Ref. Year Scope Technologies Key Contents
Li et al.[2] 2022 Socialization, healthcare, education, smart city, entertainment, real estate AI, 5G/6G, DT, MEC The authors describe the motivation and utility of using the four Metaverse pillar technologies.
Yang et al.[3] 2022 Fluid machinery pumps and fans DT, BC, AI, XR The authors investigate the application of fluid mechanical components supported by Metaverse and DT in the field of numerical simulation and fault detection.
Chang et al.[4] 2022 Education, product testing, telecommuting, production optimization, smart cities Computer vision, DT, Network, BC Three new Metaverse architectures at the edge, and technologies that help the Metaverse interact and share digital data.
Bhattacharya et al.[5] 2023 Manufacturing, internet-of-senses, industry, education, vehicle-to-everything, internet-of-bio-things Web 3.0, 6G, BC, NFTs, XR The authors discuss generalized Metaverse frameworks for modern industrial cyberspace.
Mourtzis et al.[6] 2023 Data acquisition, storage, sharing, interoperability BC, DT The authors describe the technical architecture components of BC in Metaverse Industry 5.0.
Said et al.[7] 2023 HCI XR This work illustrates the complexity of Metaverse online learning from the perspective of HCI.
Kshetri et al.[8] 2023 Economics XR, AI, 6G The global economic impact of the Industrial Metaverse.
Dong et al.[9] 2023 Task-oriented communications Semantic communication Specific requirements for the three main types of tasks in the Industrial Metaverse.
Our paper 2023 Industry 5.0 (Industrial data collection and storage, product design, operation training, system manufacturing, industrial quality control, etc.) BC, DT, 6G, XR, AI An extensive survey of the Industrial Metaverse. Particularly, For the first time, we extensively discussed the role of several key enabling technologies at different stages of industry. We comprehensively analyze the main existing problems and solutions in the Industrial Metaverse. We first emphasize more comprehensive open challenges and research directions for the characteristics of industrial scenarios.
I-A Related Works
Neal Stephenson’s novel “Snow Crash” was the first to introduce the concept of the Metaverse, a digital landscape constructed through AR technology. Since then, there has been an explosion in literature on the Metaverse, with much of it focused on its fundamental characteristics[10], empowering technologies[11], and privacy and security aspects. Some works also analyze the platform deployment and necessary components required for its application scenarios. Our research on the industrial applications of the Metaverse is comprehensive and detailed in Table II.
In terms of the empowering technologies within the Industrial Metaverse, the researchers in [6] delineate the technical architecture of blockchain (BC) in Metaverse Industrial 5.0, including its roles and limitations in data collection, storage, sharing, interoperability, and privacy protection. The work of [3] investigates the application of fluid mechanical components in numerical simulation and fault detection, supported by the Metaverse and DT. Additionally, the article by [2] expounds on the motivations and utilities of the four fundamental technologies of the Metaverse: AI, data communication, DT, and mobile edge computing (MEC). Furthermore, it describes seven crucial requirements for the current amalgamation of the Metaverse with the Internet of Things (IoT). Moreover, in [5], the authors present a cross-industry vertical Metaverse reference architecture, elucidating the underlying business logic of the contract interface as the backend, along with the front-end and back-end interaction mechanism implemented using the Web3.0 browser engine.
Figure 2: Structure of the survey.
Although some of the works do not specifically focus on industrial applications of the Metaverse, the enabling technologies involved can be directly transferred to the manufacturing industry. For instance, Dionisio et al. [10] provide a comprehensive summary of the four key characteristics of the Metaverse, namely, realism, ubiquity, interoperability, and scalability, which can be leveraged to improve manufacturing operations. Similarly, Wang et al. [12] and Yu et al. [13] analyze the potential of MEC in the Metaverse, and outline the basic components required to deploy MEC in the 6G network environment, which can be applied in various predictive maintenance scenarios in the manufacturing industry. Furthermore, Kye et al.[14] conduct a thorough analysis of the potential, as well as the constraints, of utilizing the Metaverse in educational contexts. This encompasses the domains of curriculum design and platform development, which could be effectively applied to skill training scenarios within the industry.
I-B Comparison and Our Contributions
We have noticed that the Metaverse shows great potential in the industrial sector and has gradually attracted people’s attention. Although there are numerous reviews on enabling technologies such as DT in industrial scenarios, research on Industrial Metaverse is scarce, mostly listing broad application scenarios without exploring the interconnection between enabling technologies or their roles in the architecture of Industrial Metaverse. In the work of[5], although the authors provide Metaverse architecture design for vertical industries, they do not discuss the production processes covered by enabling technologies in industrial production and analyze their specific technical support roles. The work of[15], although addressing certain industry application challenges, lacked comprehensiveness by omitting a discourse on the enabling technologies involved. Overall, at present, other research works have not formed a framework that matches the industrial production scenario, and there is a lack of in-depth combing and analysis from the perspective of industrial scenarios and technology applications. Furthermore, existing research lacks a framework that matches industrial production scenarios and lacks in-depth analysis from the perspective of industrial scenarios and technological applications. Additionally, compared to other application scenarios of the Metaverse, the industrial setting necessitates a high degree of real-time responsiveness to ensure synchronization and timely responses with the actual production environment. Any delay in the production environment could result in production line downtime or decreased efficiency. Additionally, interoperability in cross-enterprise collaboration is an indispensable feature of the Industrial Metaverse. Unlike typical commercial Metaverse scenarios, the equipment, systems, and software involved in industrial production usually originate from different vendors, utilizing diverse interface standards, communication protocols, and data formats. These elements are intertwined throughout a series of processes including design, production, testing, deployment, and feedback, each involving collaboration among various departments. Lastly, security and privacy concerns are paramount. Industrial production involves a plethora of sensitive data and confidential information such as design blueprints and industrial plans. Therefore, the Industrial Metaverse imposes stricter security requirements and must possess robust security mechanisms to ensure sensitive information is not leaked or maliciously tampered with. Hence, we believe it is necessary to comprehensively analyze and explore the application and deployment of Metaverse technology in industrial settings, and further advance related research endeavors. To this end, this paper provides a set of key contributions as follows:
  • Re-evaluate the architecture of the Industrial Metaverse, outlining its current research status and key enabling technologies, while delineating the interactions among these enabling technologies and their roles within the architecture.
  • Summarize existing research achievements on Metaverse enabling technologies and their application status in various stages of industrial production, identifying existing deficiencies for each enabling technology.
  • Analyze pressing issues in the current stage of the Industrial Metaverse, including privacy concerns, resource allocation, and platform interoperability, while examining existing solutions and shortcomings.
  • To advance industrial deployment, we summarize the current standardization efforts and open challenges in the Industrial Metaverse, and provide insights into the future development and research directions of the Industrial Metaverse, offering guidance for future work.
To the best of our knowledge, this is the first comprehensive work to organize and analyze the Industrial Metaverse and its core enabling technologies.
I-C Structure of The Survey
The sections of this paper are structured as follows. Section I presents an overview of the fundamental principles underlying the Metaverse and elucidates the rationale behind its integration into industrial contexts. Additionally, it provides an inventory of pertinent Metaverse literature reviews, comparing them to identify the innovative aspects of our current undertaking. Section II discusses around the five core enabling technologies of the Metaverse and summarizes the industrial applications and existing problems of each technology. Section III succinctly summarizes the four primary challenges encountered in the Industrial Metaverse. In addition, in Section IV we collate some of the existing standardization efforts and platforms for the Industrial Metaverse. Moving forward, Section V offers a comprehensive compilation of various unresolved hurdles prevailing in the contemporary Industrial Metaverse. Finally, our discourse culminates with a conclusive remark in Section VI. Table I indicates the list of abbreviations employed in this paper. Fig. 2 outlines the organization of the survey.
Figure 3: Industrial Metaverse Architecture. It consists of three layers: the data input layer, the enabling layer, and the industrial application layer. The enabling layer includes six components: AI, DT, BC, XR, the Metaverse management center, and the data processing system. In this context, the data input layer is composed of various industrial IoT devices. The collected multimodal industrial data is transmitted to the data processing system for preprocessing. The processed data can be stored in the BC or streamed to the AI component for analysis. The results of data analysis are used to respond to requests from the Metaverse management center, and decision outcomes are fed back to the BC. The BC performs updates on the network and DT. The DT achieves proportional replication and synchronization of virtual and physical devices and transfers parameter information to XR for 3D modeling. XR displays decision solutions through human-machine interaction devices and is applied in various industrial scenarios.
II The enabling technologies for Metaverse
This section introduces several core enabling technologies for the Industrial Metaverse, including BC, DT, XR, AI, and 6G. The primary roles of these enabling technologies in the Industrial Metaverse are as follows:
  • DT: DT provides virtual representations of physical entities and precise descriptions of products. These fundamental data are provided by industrial sensors and cameras deployed within the enterprise. Combined with AI and autonomous robotics, DT in the Industrial Metaverse is used to represent physical systems in virtual space and meet various constraints and specifications. Simulation, as a critical component of the Metaverse, is used to test vast scenarios to provide optimal modeling and decision-making.
  • AI: Integrating AI technology can enhance the modeling accuracy of DT. AI can rapidly classify and analyze large amounts of data generated from factories, supply chains, or other production processes, aiding in pattern recognition and decision-making to adjust production processes, improve production quality, and achieve the intelligence and automation of production processes.
  • XR: VR and AR serve as immersive tools providing visualization, which are more intuitive, user-friendly, and easier to interpret compared to numbers in tables or points on charts. For example, Rockwell’s development team built test scenarios on the Vuforia Studio platform, incorporating CAD files required for training tests to create wiring diagrams mapped to different product layouts and wiring schematics. Leveraging AR technology, Rockwell Automation achieved a 5% reduction in training time and extended the same training method to other production lines.
If DT, XR, and AI are prerequisites for enterprise decision-making and intelligent production in the Industrial Metaverse, then 6G and BC provide the necessary low-latency and trustworthy interaction architecture.
  • 6G: Metaverse companies use 6G networks to send and receive massive data on their Industrial Metaverse platforms. For instance, when an alert is received at the mission control center, operators can use 6G connections to remotely view and operate robots via cameras. Researchers at Nokia Bell Labs believe that compared to 5G, 6G will halve average power consumption and support a 10-fold increase in peak capacity.
  • BC: Lastly, BC, as a trustworthy distributed architecture, effectively enables trusted interactions between multiple enterprise platforms, as well as identity authentication and rights confirmation, which are indispensable in the Industrial Metaverse.
In addition, we have investigated the latest advances in theoretical research and the development of Industrial Metaverse platforms, as well as the application of each enabling technology in different phases of industrial production. In a short summary section for each technology, we briefly summarize some of the issues that still need to be addressed when integrating these technologies with the Metaverse, as well as the technical requirements in industrial scenarios.
Additionally, we have depicted the reference architecture of the Industrial Metaverse in industrial scenarios, as shown in Fig. 3. It includes three layers: the data input layer, the enabling layer, and the industrial application layer. The enabling layer consists of six components: AI, DT, BC, XR, the Metaverse management center, and the data processing system.
Firstly, the data input layer comprises various IoT devices used to collect data sources in the factory, such as cameras, sensors (e.g., acoustic/light/humidity sensors), and user-facing human-machine interaction devices like VR headsets and mobile phones. The collected industrial data includes text data, image data, video data, and other modalities. These multimodal data sources are transmitted to the data processing system for preprocessing, including data cleansing, integration, and transformation. They are stored in the BC through a BC gateway for pre-broadcasting. On the other hand, workers in the Industrial Metaverse can initiate data upload requests themselves. The BC packs and encapsulates the validated data through smart contracts, generating new blocks for continuous updates. When worker submits a data analysis request, they first need to send the request to the administrator. After preliminary authentication by the administrator, the request message is forwarded to the AI component for data analysis[16]. Meanwhile, the BC is responsible for recording transactions. The feedback from the data analysis is then used to update the DT. The DT subsequently sends it to the XR layer for 3D modeling, including modeling of production devices and environmental elements. The decisions obtained are displayed through XR headsets and serve various industrial scenarios in the industrial application layer, including product testing, skills training, and manufacturing. It is worth noting that the data used to update and adjust the DT model can come from the BC or directly from industrial equipment sensors, internal databases, or manually entered by operators. Through the DT component, the real world can be accurately replicated in digital form on multiple levels of virtual planes and can dynamically interact and synchronize with virtual devices on the virtual plane.
The communication between these layers and components is based on 6G, LoRa and WiFi communication network to provide the required QoS for the exchanged data. LoRa is a low-power wide area network technology specialized in long range communication for IoT devices. In the Industrial Metaverse, LoRa can support the connection of sensors and IoT devices to the virtual environment. It can be used to monitor and collect data in the real world and transmit this data to the Industrial Metaverse for users to analyze and utilize.
In the ensuing discourse, we shall expound upon each of these constituent strata comprising the foundational edifice.
II-A Blockchain
In the realm of the Industrial Metaverse, nodes of the Industrial Internet of Things (IIoT) amass vast quantities of sensor data in order to achieve instantaneous correspondence between the virtual and physical domains. Nevertheless, this pursuit carries with it the hazard of security and privacy breaches, as the information of these nodes is profoundly susceptible to compromise[16]. Moreover, there exists a necessity to accomplish cross-platform interoperability within the Metaverse, grounded in a specific framework.
BC, as the fundamental technology of the Metaverse, is well-suited to meet the aforementioned requirements. Firstly, it leverages identity verification and consensus mechanisms to ensure the privacy and security of users, as well as the integrity of vast amounts of industrial data, while also providing a comprehensive transaction audit trail. Secondly, due to the decentralized nature of BC, it enables collaborative production among multiple manufacturers, allowing managers to schedule and coordinate activities across multiple platforms without the need for third-party verification platforms[17]. Fig. 4 shows the vulnerabilities, advantages, and application stage of BC in Industrial Metaverse.
II-A1 Research Advances of Blockchain Enabling Industrial Metaverse
Initially, the question of safeguarding privacy in the Metaverse was explored in the financial sector, where individuals use NFTs as digital property markers to attain secure data correspondence. Each NFT is linked to an owner, and ownership records are retained on a server. Nonetheless, due to the more intricate real-time information interaction and the heavy workload associated with production tokens in the context of industrial manufacturing scenarios, it has become necessary to establish a collaborative governance framework among enterprises that is composed of more comprehensive components and architecture. To tackle this predicament, Lin et al. [18] propose a value-oriented, trusted, collaborative governance architecture as a blueprint for the deployment of the Metaverse in intelligent manufacturing scenarios. The authors provide a thorough summary of the functions and benefits of BC at various system levels, such as trustworthy data storage, secure transactions, and regulation.
It is noteworthy that computing resources, functioning as integrated asset exchanges, are also included in the domain of the Industrial Metaverse. However, the aforementioned architecture does not account for the issue of energy constraints caused by IIoT nodes. When devising an integration method between the Industrial Metaverse and BC, resource optimization and sharing should be fully taken into account. For instance, an enterprise management platform can employ various communication methods to offer computing power and encapsulate the communication steps of factory-deployed devices into a protocol. In this way, less powerful computing devices can utilize the support of stronger computing power.
Figure 4: The vulnerabilities, advantages, and application stage of BC in Industrial Metaverse.
The work presented in [19] introduces a communication paradigm for the native Metaverse, where users can access using encrypted addresses. A dedicated frequency band in the regional spectrum is utilized for the transmission of BC links in public domains. The controller can categorize access to specific resources to eliminate integration barriers between the Metaverse and inter-chain links. However, this encrypted address does not provide precise network topology information, and industrial IoT devices may possess mobility features, leading to a constantly changing topology structure that makes it impossible to determine the connection mode between nodes. Another practical concern is that profit-driven manufacturers typically consider whether the additional benefits of industrial synergy can cover participation costs. A potential solution is to design incentive mechanisms to attract companies to participate in routing-related activities. To address these two issues, the authors of [20] propose a layered BC architecture with a main chain and multiple sub-chains for sensor data analysis, combined with federated learning (FL) and cross-chain technology. Additionally, an age-based smart contract is utilized as an incentive mechanism to achieve trustworthy model aggregation and updates. Regarding the architecture design for dynamic topology scenarios, the authors of [21] adopt another approach for the architecture design, based on MEC. This solution allows MEC servers to possess additional resources for resource sharing and optimized utilization in non-trusted environments. This framework takes into account the heterogeneity of available server resources and user task requests, and utilizes a task assignment mechanism for edge resource sharing. This BC-based MEC deployment scheme can significantly enhance the utilization of computing resources while also ensuring system security.
Numerous studies have examined distinct scenarios for amalgamating the Metaverse with BC, such as architectural modeling and remote healthcare. Though these scenarios may not be relevant to our current discussion, they may potentially be customized to the industrial production domain, such as in the modeling of production devices and remote equipment control. For instance, Building Information Modeling (BIM) technology, which is utilized in 3D architectural modeling, can be exploited through parameterization methods to facilitate speedy data computation and storage for project management processes. After exploring the components of BIM, the application of the Metaverse in virtual world construction, and the latest BC research, the authors of[22] put forward a preliminary BC-based BIM data trading framework. Moreover, the author also accentuates several potential future research topics, such as immersive design experiences in the Metaverse, the evolution, and retrospection of architectural scenarios. Bhattacharya et al.[23] raise the issue of security attacks in remote surgery, which can result in the uncontrollability of the mobile robot arm. Gupta et al.[24] initially employ the hyper-ledger fabric channel in remote surgery. In subsequent work, the author further established a BC-based collaborative access control system, which comprises three parts: the master domain, the Metaverse control, and the subordinate domain, and all metadata is encapsulated in the ledger. Thus, the operator’s behavior in the Metaverse can be recorded and utilized to enhance the robot’s behavior feedback.
In general, the exploration of synergizing the Metaverse with BC is still in its nascent phase, with applications primarily focused on digital asset management, smart contracts, and decentralized identity verification to enhance the security and sustainability of the Metaverse.
II-A2 Application Stage in Industrial Production
The integration of BC in industrial applications can bring forth a multitude of benefits. One such application is in the domain of supply chain and logistics, where BC can provide a solution to the lack of traceability and transparency in traditional supply chain management. By enabling the tracking of the entire life cycle of a product, from raw materials to finished goods, BC can ensure safe and convenient management of commodity flow[25].
In industrial data collection, the non-repeatability of data is crucial for accurate data analysis and effective business decisions. With BC, the hash value of each data block can be calculated and compared with the previous block’s hash value to ensure that the collected data is not duplicated, thereby improving the reliability of data collection.
In terms of data storage, the massive amount of multimodal data generated during industrial production processes can quickly exhaust physical storage capabilities[6]. Overreliance on centralized storage systems in the Metaverse can result in a single point of failure, leading to significant economic loss. BC’s decentralized paradigm can effectively solve the problem of single point of failure while preventing data tampering through the continuous creation of new blocks.
Furthermore, BC can promote cross-enterprise production collaboration and platform interoperability by realizing cross-chain protocols and facilitating cross-chain transactions and value flow across different devices[26]. At the enterprise level, it can enable collaboration among suppliers, manufacturers, and end-users, as well as cross-chain governance among different enterprises during the production stage.
II-A3 Summary
BC can provide efficient management for the Industrial Metaverse platform through mechanisms such as non-tampering, traceability, collective maintenance, and openness and transparency[27]. However, in a BC network, each participating node must verify transactions, which can lead to latency issues due to network transmission, disk writing, and other operations. Although several studies have been proposed to optimize the network topology, design new consensus algorithms, and adopt a hierarchical structure to alleviate this problem, these solutions are limited by the size of the network and the number of nodes. Therefore, balancing decentralization, security, and scalability under the Industrial Metaverse architecture is an urgent problem that needs to be addressed for sustainable enterprise development.
II-B Digital Twin
DT can produce a virtual replica of a physical factory by gathering and analyzing production data and equipment parameters in the actual facility. With this information, DT can perform behavior predictions of physical objects or machine states. By accurately mirroring the real world in the virtual world, DT can resolve complex issues that would otherwise prove challenging to solve.
Hence, to develop an Industrial Metaverse platform, DT must fulfill three vital requisites: the capability to simulate industrial entities, the capacity to dynamically procure data from the live environment, and the proficiency to synchronize the data in real-time. Fig. 5 shows the vulnerabilities, advantages, and application stage of DT in Industrial Metaverse.
II-B1 Research Advances of DT Enabling Industrial Metaverse
DT made its debut in the industrial sphere in 2003, when Michael Grieves proposed its implementation in product lifecycle (PLC) management. By utilizing production data and digital models, simulated processes became possible. As the IIoT and DT evolved, the primary focus shifted towards using DT to reduce industrial sites and enhance product evaluation through assistive technologies.
However, some researchers argue that DT tends to overlook crucial action details and treat operators and small devices as mass points in simulations, affecting object modeling. To overcome this challenge, Zhou et al.[28] design a small object detection framework that fuses multidimensional sensor data, treating devices, products, and operators as fundamental environmental parameters. They implemented multi-level feature fusion based on hybrid neural networks created with MobileNetv2, YOLOv4, and Openpose. To effectively reduce the false detection rate, Wu et al.[29] utilize multi-modal fusion data as a data source, Cyber-Manufacturing Systems (CMS) interface for numerical insertion and visualization in DT for product surface defect detection tasks. Finally, they employed morphological alignment algorithms. Numerous studies have revealed that improving the simulation capability of DT requires integrating and optimizing AI algorithms.
Furthermore, given that DT necessitates substantial bandwidth and computing resources, it inherently demands real-time synchronization with stringent reliability and low latency prerequisites[30]. As a result, achieving real-time synchronization for DT entails stringent reliability and low latency requirements. Therefore, an important research topic is the deep integration of DT into industrial platforms. This includes designing a shared architecture for digital factories and developing scheduling algorithms to optimize resources. A novel approach is to utilize DT in conjunction with edge intelligence to facilitate immersive Metaverse connectivity. By optimizing resource allocation, ultra-reliable and low-latency communication (URLLC) can be realized[31]. The authors in[32] explore the bi-directional dependency of DT with MEC and the interaction of DT-assisted edge computing with DT to avoid Metaverse service disruptions. Dai et al.[33] integrate lightweight DT with MEC to reduce cloud load, increase the number of service requests for MEC, and significantly reduce the time complexity of the algorithm by introducing Merkle trees. Although there have been numerous studies on industrial DTs for handling edge selection and task offloading, these schemes often overlook the data storage limitations of the Metaverse platform. In fact, the fundamental components of the Industrial Metaverse can be simply divided into virtual service providers (VSPs) and individual users and enterprises. However, not all physical assets can be deployed on the platform to capture device status information in real-time to ensure full interoperability. Han et al.[34] propose a dynamic layering of platforms through IoT devices, where VSPs maintain business profitability by increasing synchronization strength. The IoT devices are responsible for transmitting data to independent VSPs. The proposed architecture achieves optimal synchronization through optimal control theory. Zhang et al.[35] investigate the problem of multi-base station channel resource allocation in a Metaverse DT, using an end-to-end cloud three-layer framework. They design a reverse auction mechanism to solve the resource allocation and pricing problem. Furthermore, the authors in [36] consider the joint design of communication, computation, and storage resources to formulate the problem of minimizing latency by optimizing edge task offloading and arithmetic overhead. This improves the Quality of Experience of DT in the Metaverse.
Figure 5: The vulnerabilities, advantages, and application stage of DT in Industrial Metaverse.
DT enables the digitization of physical objects, while XR technology provides the medium for interaction between the two. Therefore, in addition to DT itself, a part of the works investigate the synergy between DT and XR technologies in industrial scenarios. For instance, Yang et al.[37] propose a framework for smart crane applications based on DT and XR to improve program development efficiency and Human-computer Interaction (HCI) system usability. Coupry et al.[38] investigate options for XR technologies to improve BIM and highlight the challenges of implementing BIM-based systems. Jerov et al.[39] investigate the prospects of DT and XR applications in control systems, where they develop a DT system for XR with basic interaction mechanisms and connect it to real production scenarios. While some of the above works are limited to conceptual design and specific use cases, a collaborative approach is needed to build a system that integrates DT and XR effectively.
Tu et al.[40] propose a TwinXR approach that effectively merges information management DT with knowledge-based XR technology. By leveraging IoT devices available in the factory, XR programs can be swiftly developed and instantiated, while the development process is greatly facilitated through the use of publicly available software packages. The authors effectively demonstrate the remarkable compatibility and versatility of this approach by showcasing its successful application to cranes and robot arms in smart factories. Nonetheless, the authors duly acknowledge that the adoption of TwinXR methods can be prohibitively expensive when the machinery is stationed at a specific, static position, rendering cost a critical limiting factor in real production settings.
It is noteworthy to mention that, in addition to the collaborative architecture research, the real-time simulation characteristic of DT allows for monitoring the dynamic properties of physical portals for detecting various network attacks. Gupta et al.[41] suggest synchronizing the device-generated traffic through the gateway router, utilizing Support Vector Machines (SVM) for achieving regular synchronization with DT. This framework attains an accuracy rate of over 93% in detecting malicious traffic.
II-B2 Application Stage in Industrial Production
Product design stage: During the product conception stage, companies utilize DT to test the production environment, constructing a virtual model to improve production by adjusting parameters in order to reduce design defect rates and production costs, while meeting preset requirements. This process entails continuous iterative optimization of both digital and physical models, interconnecting through data flow[42].
Supply chain and Operations & Maintenance (O&M): DT can optimize the production process in remote equipment O&M and supply chain management. In equipment operation and maintenance, DT provides real-time data on equipment performance, including its current condition, remaining useful life, and any potential problems or failures that may occur. In the supply chain management segment, DT creates a digital copy of the supply chain network, enabling managers to monitor and optimize their operations in real-time.
Post-commercialization services: Furthermore, the Industrial Metaverse platform integrated with DT can be used to manage the traditional industrial service phase, providing DT information on the product after it has been finished and put into the market. Both manufacturers and technical service providers can access product status information on the cloud platform, which can be used to help engineers and designers improve customer experience through customized services.
II-B3 Summary
To put it succinctly, DT enables the integration of physical and digital worlds through real-time interaction and data analysis, providing a new perspective for addressing industrial challenges[43]. By combining explicit and tacit knowledge, real-time operational data can be incorporated into traditional industrial models to build DT that aid in industrial management. However, further research is needed to consider constraints within the Metaverse[44], such as security and storage, as well as to explore integration with other enabling technologies, as current efforts in this area are still in their infancy.
II-C 6G Communication Network
The Industrial Metaverse is typically data-driven and sensitive to perception latency at the physical layer. However, sensor data at the physical layer in industrial scenarios is often complex, with characteristics such as high volume, multiscale, and high noise. Additionally, the supporting technologies involved in the Industrial Metaverse often impose high demands on communication bandwidth. For example, NFTs consume significant bandwidth during digital asset validation, and the verification of digital tag sets in BC, as well as the bandwidth requirements for transactions involving different entities[45]. Furthermore, VR systems require a bit rate of 1 Gbps to provide qualified video streams visually, and real-time feedback for remote operations also necessitates high reliability and high data connection rates.
5G’s URLLC are both unable to meet the aforementioned requirements. The next-generation wireless communication technology represented by 6G can provide high throughput and low-latency guarantees for the interaction between the physical and virtual worlds’ upper and lower-layer links. This is achieved through fast transaction processing to achieve efficient management. In addition to throughput, the goal of the Industrial Metaverse is to ensure high concurrency of devices in space and time, allowing all devices and users on the platform to access and leave anytime, anywhere. This increases the demand for seamless and comprehensive connectivity provided by 6G. In this regard, 6G can provide auxiliary connections through coordination with IoT devices and achieve global coverage, while its intrinsic security technology can be used to defend against unknown security threats. Fig. 6 shows the vulnerabilities, advantages, and application stage of 6G in Industrial Metaverse.
II-C1 Research Advances of 6G Enabling Industrial Metaverse
Most existing works are optimized for 6G networks and integration with the Metaverse, and metric focus varies for different application scenarios. The performance metrics in the Industrial Metaverse include throughput, latency, reliability, Age of Information, and energy efficiency[46], among which all metrics’ weights, except latency and reliability, are deeply connected and affect joint architecture design. In this regard, Cao et al.[47] discuss several metrics’ coupling relationship in-depth and argues that communication delay, reliability, and AoI are the three most crucial metrics in an IIoT environment with heterogeneous traffic. The authors collect data and control heterogeneous traffic from the short packet transmission optimization perspective. However, their specific task division is not given, and transmission jitter’s performance loss is not negligible in the actual industrial environment.
Another approach is based on overall architecture design. MEC, an essential auxiliary mechanism of the 6G network, can significantly reduce request delay due to its proximity to the source. The devices in the factory, following the edge computing hierarchy, can include sensors, IIoT devices, and servers[12]. Sensors collect industrial data, and various IIoT device states are responsible for receiving and then converge to the cloud for unified processing. In a typical scenario, some lower-level operations in industrial production can be decentralized to the lightweight mobile end, and the administrator can control the device remotely. Various 6G mobile edge technologies exist for application to the Metaverse, including wireless communication technologies (heterogeneous radio, intelligent reflective surfaces, non-orthogonal multiple access, etc.) and new communication technologies (semantic communication, holographic-type communication, tactile communication, etc.)[13].
Aslam et al.[48] present a layered Metaverse design from the perspective of layers, with different layers connected by a data exchange layer. They then demonstrate the autonomous driving remote control case and finally discuss the challenges and propose solutions for 6G and Metaverse integration, including how to deal with wireless communication security threats, submillimeter beamforming loss, and channel interference.
Since 6G mobile authorized devices in the same production environment need to share computational memory and frequencies, unreasonable resource allocation can easily lead to network congestion, which we will cover in detail in the next section.
Figure 6: The vulnerabilities, advantages, and application stage of 6G in Industrial Metaverse.
II-C2 Application Stage in Industrial Production
Computer Vision Scenarios with Multiple Accesses: With the advent of 6G-enabled industrial IoT devices, computer vision scenarios can be significantly enhanced with seamless connections and high data transfer rates, enabling better image quality in the Metaverse. Additionally, MEC resources in a virtual 6G network can leverage deep learning techniques to process heterogeneous devices in the enterprise[45].
Product Decision and Service Phase: The product decision and service phase can also benefit from 6G, as it facilitates computation-centric communication by exchanging communication resources to ensure the required quality of service for adequate computational accuracy. This, in turn, improves decision-making efficiency and enables further automation. Moreover, 6G can simplify the supply chain network, leading to agile and adaptive service provisioning for Metaverse applications.
II-C3 Summary
The network heterogeneity of each Metaverse subplatform of manufacturers presents a significant challenge for the development of a unified communication network, requiring a common network access and control protocol. Additionally, plants that deploy 6G devices must ensure the compatibility and operability of existing network-enabled devices, which must be fine-tuned to each individual subsystem. Currently, there is a lack of available 6G-enabled IIoT architectures for deployment on Industrial Metaverse platforms, as well as a lack of corresponding standards and specifications. Although relevant reference architectures have been proposed[49], the hierarchical relationships are oversimplified and require further empirical testing, and the theoretical aspects remain unexplored.
II-D Extended Reality
In manufacturing scenarios, there are instances wherein it becomes imperative to employ diminutive sensors or IoT devices at the distant terminus, which typically lack interfaces conducive to Metaverse user engagement. XR, including VR, AR, and mixed reality (MR), serves as the interface between the tangible realm and the ethereal domain of the Metaverse, endowing users with a myriad of multifaceted interaction modalities[50]. Within this realm, VR unveils itself as a veritable medium for sensorial emulation, employing head-mounted apparatuses to engender immersive experiences. It facilitates the visualization of diverse design iterations, verifies procedural efficacy, and bestows captivating verisimilitude. On the other hand, AR seamlessly integrates virtual components, such as textual and pictorial elements, into the fabric of reality through handheld devices like smartphones and tablets. By availing themselves of this technology, operators are endowed with instantaneous access to real-time equipment status updates, virtual cues, and instructional guidelines. Furthermore, MR transcends conventional overlays of virtual content upon tangible surroundings, instead forging a harmonious convergence between the virtual and the corporeal, facilitated by perceptive and interpretative faculties attuned to the veritable realm.
In broad terms, constrained by diverse factors in industrial scenarios, comparatively substantial apparatuses like robotic arms, for instance, manifest greater aptitude for remote manipulation by technicians, whereas XR can function as an ad hoc controller. Furthermore, XR technology has the capability to furnish a fusion of tangible and virtual domains for user-machine interplay, thereby lending itself to deployment within industrial production scenarios for the purposes of operational guidance, equipment maintenance training, and product modeling and design. Fig. 7 shows the vulnerabilities, advantages, and application stage of XR in Industrial Metaverse.
II-D1 Research Advances of XR Enabling Industrial Metaverse
In a certain context, XR constructs an adaptable and fluid realm for visualizing virtual content within the realm of MR. However, a primary concern arises regarding the imperative of upholding the sustainable connectivity attributes of content within the physical realm, as virtual contrivances in the Industrial Metaverse persistently expand. Should an information disjunction arise between the Metaverse and the physical space, coupled with a dearth of consistent communication, it can potentially engender cognitive overload for the user[51]. Put simply, when operators engage in interactions within the virtual plant, they are typically unable to access and manipulate the genuine physical apparatus. Consequently, it is crucial to address these disconnections within the Metaverse in a scientifically comprehensive manner, as failure to do so can easily impinge upon critical industrial scenarios.
An intriguing concept in this realm is the utilization of an extended Metaverse agent approach based on the MiRAs cube theory[52]. This approach serves the purpose of dynamically governing virtual device control and hologram modeling, while facilitating interactions with other holograms and operators to manifest within the physical realm. Building upon this notion, Rincon et al.[53] leverage agents founded on the meta-model (MAM5) and JaCalIVE frameworks to mitigate discrepancies between simulated and real systems. Intelligent resource artifacts are incorporated, enabling seamless access to the physical world for devices. The dynamic essence of agents is further explored by Croatti et al.[54], whose architecture allows for agents to be dynamically altered at runtime, employing diverse heterogeneous agent programming techniques to modify agents within AR. In addition to system modeling, certain researchers have embraced agent-based approaches to automate XR quality testing and support experience evaluation by adopting user roles[55]. Braud et al.[56] construct an XR architecture from the operating system level, which perceives the physical world as a shared resource among applications. By implementing meticulous control mechanisms through several components such as environment understanding, specialized chip drivers, networking, interoperability, and various other interconnected elements, they achieve enhanced integration of content with the physical-digital world, thereby attaining superior performance. Drawing upon the extended Metaverse architecture, Guan et al.[51] conduct an in-depth analysis of interface design consistency within the Metaverse. They attribute inconsistencies to signal noise and propose the utilization of ”noise reduction” techniques to mitigate disparities between the Metaverse and the physical world. This approach aids in facilitating the development of interoperable interfaces. The authors further delve into the realm of seamless integration by exploring the presence of interoperable agents. They showcase the effectiveness of this approach through two early design projects, presenting interactive and tangible Metaverse applications that can be deployed across both physical and virtual realities. As interfaces of this nature continue to be researched and developed, the integration of suitable filters and controls becomes crucial in ensuring the sustained continuity of the Extended Metaverse. This integration holds the potential to serve as a valuable research direction in the long-term future.
Another intriguing concept in interactive device interfaces is the utilization of a convergent technological paradigm known as XR-IoT (XRI) systems, which combine XR and IoT. These XRI systems capitalize on the complementary attributes of both domains, leveraging the interconnectivity of environmental entities facilitated by IoT, while XR empowers information-rich interactive interfaces and serves as a platform for HCI[57]. The evaluation of interface design in XRI systems is initially proposed by Dnser et al.[58], who delineat the original design intent of XRI and presented a comprehensive evaluation framework from the perspective of interface design, with usability at its core. Building upon this work, Tsang et al.[59] provide a more exhaustive taxonomic approach and a QoE reference standard for the development and assessment of XRI systems. In the context of real-world industrial scenarios, Morris et al.[60] envision an immersive, multi-user, and multi-agent-driven XRI-based intelligent workstation system. This system incorporates haptic feedback-equipped desks, cameras, and other devices to monitor personnel status, while physical laptops run machine learning models to determine personnel conditions. Taking it a step further, Oppermann et al.[61] introduce a 5G XR toolbox tailored for remote maintenance scenarios in industrial equipment. This toolbox facilitates MR and VR views by establishing digital connections between field equipment and personnel. Operating on actual machines within real production lines, the system utilizes authentic CAD data to create a realistic prototype environment.
Another intriguing perspective to explore is whether the utilization of AR/VR devices can have adverse effects on workload, considering that workload predominantly reflects the input costs in production tasks. Kantowitz et al.[62] demonstrate that efficiency tends to decline as task complexity increases during HCI, even within the maximum workload capacity. To put it simply, it raises the question of whether operators need to exert additional cognitive effort to execute tasks through VR interfaces that could otherwise be performed in the real world. In this context, Xi et al.[63] conduct an investigation on the workload disparities of XR across six distinct dimensions. Their finely detailed analysis reveal that AR has a more pronounced impact on workload compared to VR, while the combination of both technologies do not escalate task difficulty. Future research endeavors can delve into exploring the mechanisms and boundary conditions of these different dimensions, enabling XR designers to develop more scientifically informed HCI systems.
Figure 7: The vulnerabilities, advantages, and application stage of XR in Industrial Metaverse.
II-D2 Application Stage in Industrial Production
Product Recognition and Modeling: Object detection is widely applied in XR and is an essential task for achieving the Metaverse. For example, computer vision techniques are used for real-time recognition and overlaying relevant information on the display screen, such as product specifications, usage instructions, and maintenance methods. It can also be applied to the modeling and visualization of industrial products, where real-time visual feedback helps improve the efficiency of product design, assembly, and maintenance.
Device Pose Recognition: In XR, it is often necessary to observe and recognize the actions of devices like robots in a 3D immersive factory environment and provide feedback for specific actions. For instance, on an assembly line, AR technology can be used to detect the pose of devices or components in real time, including rotation angles, positions, and sizes, thus assisting operators in correct assembly and adjustment.
II-D3 Summary
The Industrial Metaverse is a collaborative virtual space that enables enterprise employees to interact and work together in a digital environment. By manifesting as tangible virtual images within this space, operators can engage in virtual production and collaboration that runs parallel to the real world, thereby enabling digital transformation. This immersive technology is poised to shape the new form of the Industrial Internet. Additionally, AR and MR technologies have the potential to revolutionize the operation of the physical world within factories. As digital entities transition from a purely virtual environment to a physical one, further industrial design and technological considerations are necessary to ensure the success of digital transformation.
Here’s the table without the blue font color commands:
TABLE III: Research on enabling technologies in the Industrial Metaverse.
Tech. Ref. Year Scenes Objectives Methods Limitations
BC [18] 2022 Data collection to storage. Ensure credible execution of Industrial Metaverse. A BC-based trusted collaborative governance system for Metaverse smart manufacturing scenarios. Lack of simulation.
[20] 2022 Cross-enterprise collaboration. Improving privacy protection in the Industrial Metaverse. Combining joint learning and cross-chain techniques to design a hierarchical BC architecture with main chain and sub-chains. Lack of mechanisms to prevent malicious local devices.
[19] 2022 Access management. Responding to Metaverse and decentralized Internet initiatives. Dedicated frequency bands in regional spectral bands for link transmission. Lack of simulation.
[21] 2022 Cross-enterprise collaboration. Optimizing resource sharing and utilization. A practical BC-based MEC platform is proposed for resource sharing and optimal utilization to accomplish requested offloading tasks. No detailed assessment of the proposed system is provided.
DT [36] 2022 Cross-enterprise collaboration. Joint design of communication, computing and storage resources. The latency minimization problem is formulated by optimizing edge task offloading, arithmetic overhead, and thus improving the QoE of DT in the Metaverse. The bias is assumed to be fixed.
[35] 2022 Base station data receiving. Ensuring the quality of VSP’s services. The authors investigate the multi-base station channel resource allocation problem in a Metaverse DT with an end-to-end cloud three-layer framework. Storage resources are not considered.
[29] 2022 Surface defect detection. DT solutions for small surface defect inspection tasks. Integration of multimodal data. Algorithms struggle to handle massive data scenarios.
[33] 2022 Base station data receiving. Reduce cloud load with lightweight DT integration with MEC. A Service Placement Algorithm Based on Merkle Tree. The baseline algorithm is not enough.
[37] 2022 Crane application. Improved efficiency of XR application development. The architecture includes a perception layer and ROS-based simulation model, and a Unity-based service layer to create XR-based digital applications. It may not be possible to apply it to other areas.
[40] 2023 Supply chain and O&M. Promoting synergy between DT and XR. The TwinXR approach is proposed, which combines information management DT and knowledge-based XR. Feature mapping is entirely dependent on shared data structure.
6G [4] 2022 Computer vision. Organic integration of edge AI and Metaverse with 6G support. Three novel edge Metaverse architectures are introduced that use 6G-enabled edge AI to address resource and computational limitations. Missing network latency and privacy issues with 6G.
[64] 2022 Product decision and service phase. Allows Metaverse mobile users to download virtual world scenes in real time. An environment with multiple cellular stations is designed where the task of downloading graphics would be handed over between cellular stations. The role of AI for 6G is not addressed.
XR [54] 2018 Hologram modeling. Developing agent-based AR/MR systems. Support for modifying agents in AR using different heterogeneous agent programming techniques. Lack of integration with standards.
[65] 2022 Interaction & Assistant. Adapting to MTMM workload requirements for device heterogeneity. Various features of real-time MMMT ML workloads are discussed, and XRB ENCH and three representative executions are introduced. No consideration of cascading pipeline complexity.
[66] 2022 None. Understanding the relationship in XR’s perception of user privacy. The article analyzes the gap between user privacy perceptions and their specific behaviors across XR. Omission of audit and transparency.
[56] 2022 Automated quality tests. Operating system level XR architecture. Get better performance by enforcing fine-grained control. No mention of actual use cases for XROS.
[67] 2022 Workload metrics. Correlation of AR to overall workload. NASA-TLX for measuring subjective workloads. Subjective experiment.
AI [68] 2022 Multi-Team collaboration. Lightweight Metaverse framework to support language and semantic. Summarizes common AI technologies and application development templates, as well as common functional modules and interfaces. Lack of quantitative assessment.
[69] 2022 Product design. Improve 3D modeling efficiency. AI-assisted VR modeling. Lack of baseline comparison.
[70] 2022 Product design. Memory interfaces are too restrictive in terms of bandwidth. Higher XR resolution can be achieved by stacking multiple layers of memory using 3D stacking. Lack of practical examples for deployment.
[71] 2022 Digit recognition. Mitigating security risks in the Metaverse. Arming smart contracts with intelligence and using AI algorithms for prediction. Missing energy overhead.
[72] 2022 Product performance analysis. Improving the reliability of numerical modeling. Four ML methods are used to fabricate advanced DTs of 42SiCr alloy considering all their uncertainties and nonlinearities. The industrial scenarios involved are relatively homogenous.
II-E Artificial Intelligence
As the foundational cornerstone of the Industrial Metaverse, AI obtains its input data through an array of sensor devices situated within the factory. This data, once subjected to meticulous preprocessing procedures, serves a multitude of industrial production processes, encompassing the realms of image and audio recognition, astute decision-making support, and production planning, among others[73]. Concurrently, AI remains attuned to user interactions transpiring within the virtual realm. Furthermore, beyond its inherent industrial applicative worth, AI possesses the potential to harmoniously unite with other pivotal enabling technologies, thereby bestowing an enhanced HCI experience. All in all, the pivotal role of AI technology lies in its provision of essential implementation reliability and exceptional performance standards to propel the Metaverse forward.
II-E1 Research Advances of Artificial Intelligence Enabling Industrial Metaverse
At this juncture, the research endeavors surrounding AI in the Industrial Metaverse converge upon two primary focal points. Firstly, the synergy between AI and other enabling technologies takes center stage, encompassing algorithm optimization for VR devices, as well as deployment paradigms stemming from the amalgamation of BC and MEC. Such combinations give rise to distributed AI and edge intelligence, which find practical application within the Industrial Metaverse. Secondly, dedicated technical research pertaining to industrial AI itself unfolds, including domains such as natural language processing, speech recognition, and the notable emergence of generative AI models, typified by chatGPT, which exhibit broad applicability in domains like intelligent Q&A and assisted decision-making.
First, under AI-assisted mechanisms, XR includes technologies such as computer vision and XR computing, which can immediately capture environmental data within the field of view and convey real-time processing results to users through display devices. On one hand, research has been conducted on adaptive architectures for XR devices. Wong et al.[69] devise a system called GetWild, where the main network in the AI stage is CGAN. This stage allows users to capture images using VR devices as input to create rough 3D objects and terrain, and further perform editing operations. This technology can be effectively applied in the industrial product design process. However, this massive 3D modeling can pose resource constraints for small edge devices such as VR goggles. Even with optimizers for XR device workloads, latency issues caused by high memory requirements still exist. To address this, Yang et al.[70] employ a 3D stacking neural network accelerator architecture to evaluate energy and latency improvements. The simulator provides estimates of various hardware metrics, such as execution time (latency), energy, input/output flow, and resource utilization of prototype ML accelerators for running AR/VR models. Based on the analytical model of the simulator, operators are extracted, and expected performance indicators are calculated based on the energy table of the accelerator architecture to guide architecture design decisions for different ML models. Subsequent work can further expand the memory capacity in the X, Y, and Z directions to identify points of diminishing returns and consider the practical deployment and implementation of the platform. In addition, for VR performance estimation, Yang et al.[74] contribute a dataset (VRQ-TJU) that can be used for VR quality assessment and introduced an end-to-end 3D CNN for video quality assessment. The 3D CNN takes VR differential video patches as input and considers information between different frames. Subsequent work can analyze various features rather than just spatial distribution changes. In addition to auxiliary architecture optimization and quality assessment, AI is widely used for VR image denoising, super-resolution generation, etc., to provide operators with smaller granularity of operations and clearer interactive interfaces[75],[76].
Another trending direction entails AI-assisted biometrics for human pose recognition in complex environments and user device verification within the Metaverse’s realm. The former encompasses motion recognition and body part tracking facilitated through the synergy of controllers and motion sensing devices. Such applications commonly manifest in player interaction scenarios within gaming contexts, but fall outside the purview of this discourse. The latter, primarily rooted in user privacy protection, will be expounded upon in the ensuing section.
In conjunction with the integration of XR, the Metaverse architecture empowered by BC, Web 3.0, and MEC has engendered the emergence of Decentralized AI (DeAI)[77]. In contrast to Centralized AI (CeAI), which embodies a static, centralized system leading to limited flexibility and scalability, DeAI possesses the capability to decompose tasks assigned by central departments into subtasks, subsequently allocating them to the next tier of agents for processing. Through collaborative efforts, these agents ultimately converge to formulate the final solution. While intricate AI models often entail substantial computational overhead, rendering real-time execution on edge devices within the Metaverse a challenge, lightweight AI processing solutions have matured. These encompass techniques such as parameter pruning, quantization, low-precision approximation, and knowledge distillation. Lee et al.[78] have successfully applied a lightweight LSTM model in Head-mounted Displays (HMDs). The network takes a set of 10 data points sampled from motion data at intervals of 5 timestamps as input. Subsequently, online learning techniques can be applied for personalization to enhance accuracy. By virtue of lightweight edge deployment, DeAI enables the realization of intelligent device agents at the edge level. Nodes within the edge layer undertake preliminary solutions through local reasoning and subsequently coordinate and communicate with neighboring nodes, thereby facilitating resource sharing. This paradigm presently represents the optimal approach for the Industrial Metaverse.
Another facet of AI-enabling technology revolves around the advancements within AI technology itself. With the escalating popularity of generative AI models such as AI generated content (AIGC), the capacity of AI to generate content has emerged as a new frontier within the Metaverse. Consider an employee equipped with an AR device tasked with arranging equipment in a virtual environment. In this scenario, the employee may require a catalog of pre-ordered equipment, prompting the input of relevant details into an AIGC model. The model then generates equipment entities based on the given context and scenario. Notably, AIGC within the Metaverse should meet two essential criteria: first, it should support device access at any time and any location; second, it should maximize user utility within the Metaverse while satisfying user needs. Nevertheless, due to variations in task suitability among AIGCs and the non-uniform distribution of computing power across edge servers, a resource scheduling scheme that ensures reasonable allocation becomes imperative. Addressing this concern, Du et al.[79] propose an optimal decision algorithm for AIGC based on a diffusion model, used to generate ASP selection decisions. The system collects AIGC task requirements from users, where user submissions include text prompts describing the desired content. Edge servers process and schedule these requests, ensuring efficient resource utilization and optimal system performance by understanding the specific requirements of each task, including computational complexity and output quality. After collecting user requirements, the system conducts semantic analysis through a knowledge graph to identify similarities and differences between user prompts. This approach provides an effective solution for deploying AIGC services in the Metaverse. Similarly focusing on edge deployment for AIGC, Xu et al.[80] synthesize the trade-offs between reasoning latency, resource consumption, and accuracy by managing pretrained foundation models (PFMs) on edge servers using the minimal context algorithm, reducing the overall system cost. This work provides a formal mathematical expression for basic model caching and inference problems and the minimal context algorithm. Overall, the emergence of AIGC illuminates new avenues of development within the Metaverse, enhancing content diversity and alleviating repetitive production labor. However, it also gives rise to novel challenges, such as the regulation of generative content and the screening of potentially explosive content[81].
AI is poised to enable intelligent networking, immersive digital worlds, inclusive user interfaces, accurate avatar creation, multilingual accessibility, and many other features within the concept of the Metaverse, which is inherently user-centric[82]. The more precise and refined the AI models are, the better the overall experience can be for users. Multilingual accessibility and automatic translation are also critical components that facilitate cross-border enterprise access to the Metaverse[83]. Nevertheless, interpretable AI is essential to enhancing brain mapping accuracy, as it enables explanations for trained models and enhances the interpretability of AI black box operations. In recent years, a variety of interpretable AI-based approaches have been introduced to improve control and monitoring of unwanted or adverse effects, such as biased decision making. Interpretable AI utilizes semantic representations to uncover relationships between different object coverages in images, thereby enhancing the interpretability of deep learning solutions. Additionally, interpretable AI can be utilized for product design and marketing, helping to analyze competitor performance and identify opportunities to outperform them[84]. By leveraging customer-generated reviews, interpretable AI can extract competitive factors for products and effectively reflect customer opinions.
II-E2 Summary
AI technologies in the Industrial Metaverse generally require meeting requirements such as real-time capability[85], scalability[86], and interpretability[87]. However, there is still a significant research gap in interpretability. Many AI-assisted services and applications in the Metaverse rely on AI agents that are driven by black-box ML models, lacking explanation and interpretability capabilities. Developers, virtual world designers, and users of the Industrial Metaverse cannot understand the AI decision-making process from a mechanistic perspective. The lack of interpretability may lead to potential risks, as even a 1% error rate in decision-making can be unacceptable in industrial manufacturing. This impact not only affects economic benefits but also safety concerns.
III Existing issues in Industrial Metaverse
In this section, we will delve into the crucial issues that currently exist in the Industrial Metaverse. With the rising complexity of connections and interactions between systems and devices, security threats and vulnerabilities have become a more pressing concern. Furthermore, the vast number of sensors, devices, and data involved in the Industrial Metaverse require efficient data management to prevent privacy breaches during data transmission[88]. Given that most enterprises prioritize profit maximization and face resource constraints, scientific resource optimization is necessary to reduce costs. We will elaborate on these challenges in the following sections. Fig. 8 illustrates the existing issues in Industrial Metaverse.
Figure 8: Existing issues in Industrial Metaverse. Including security and privacy, resource allocation, and interoperability. In terms of security, it can be discussed from the perspectives of sensors, platforms, and enabling technologies themselves. Resource allocation includes two aspects: computing resource allocation and storage resource allocation.
III-A Security and Confidentiality
In the preceding segment, we presented an overview of how the Industrial Metaverse has attained extraordinary immersive encounters through the facilitation of cutting-edge technologies such as XR, DT, and AI, thereby establishing a novel paradigm for industrial administration. However, at the same time, it has also created unprecedented security threats. From data collection to the final decision-making stage, security and privacy challenges pervade the entire production cycle of the Industrial Metaverse. For example, the requirements for data integrity and availability, which are based on industrial data itself (i.e., ensuring that data flows from sensor collection endpoints to DT models without being maliciously modified or destroyed), as well as the requirements for authentication/authorization and traceability for employees (i.e., strict verification mechanisms for device operations and data access to ensure that only legitimate users can access them).
Although in the previous section, we mentioned the use of BC technology for anonymizing data and user identities through techniques such as identity verification and smart contracts, this alone cannot solve all security issues. In the following, we will organize existing solutions to security and privacy issues existing in various stages of the Industrial Metaverse based on the architecture diagram provided in Fig. 2 This includes data sources collected at the sensor end, as well as data and network security threats during subsequent cross-component transmission processes. We will also discuss security issues within components such as AI, BC, and XR, and further discuss possible research directions.
III-A1 Sensor-side Industrial Data Source
To enhance the immersion of the Metaverse and achieve more precise industrial decision models, industrial sensors play a pivotal role in collecting extensive data for the platform. These data sources encompass not only initial training samples and physical parameters used to construct the virtual space but also user information.
When it comes to industrial source data, the emphasis lies in safeguarding the data itself. For instance, redundant information in visual content, which does not contribute to the training of models but is indiscriminately collected by sensors, can be obfuscated by overlaying a mask to mitigate potential privacy breaches. Validated operators can have their virtual images in the Metaverse replaced with generic mappings. Another approach involves employing black and white list mechanisms, where only specific information is retained, and any data outside the whitelist is immediately discarded and inaccessible. Wu et al.[89] propose a data misuse prevention mechanism called DAPter, wherein input data undergoes lightweight generator transformations to remove unnecessary information with minimal additional overhead. The blacklisting mechanism is commonly utilized for safeguarding user facial information. In this method, faces are decomposed into vector signals, comprising identity and attribute features. Subsequently, an anonymization model is generated to integrate and reconstruct the vectors based on the need for further processing[90]. This approach can also be employed for classifying features of industrial devices, such as device models and metric parameters. Additionally, traditional data encryption and storage schemes used in cloud-side end systems can be applied to Industrial Metaverse platforms. These include classical symmetric and asymmetric encryption techniques, cloud storage encryption, disk encryption, and more.
Apart from the issue of data leakage, the quality of data sources is also classified as a data security concern. On one hand, malicious users may generate low-quality or even false data to disrupt virtual world project construction. On the other hand, profit-oriented participants may lower data verification standards to control costs, leading to misleading data analysis and model training. The first case can be addressed through identification of malicious users and tackling data-related attacks. Common data attacks include data injection attacks and data tampering attacks. Data injection attacks involve perpetrators injecting false data into industrial systems through illicit means to mislead and disrupt system operations. Countermeasures against such attacks include implementing detection and verification mechanisms, as well as access control. For example, Pang et al.[91] employ a Kalman anomaly detector based on residuals to securely control networked systems, while Guo et al.[92] use a random coding scheme to detect residuals, transforming the design problem of coded signals into a constrained nonconvex optimization problem. Access control, which restricts data modification to authorized personnel only, is widely employed in industrial settings. Existing schemes include access control architectures based on BC[93], cryptographic storage methods[94], and more. These access control schemes can also be applied to combat data tampering attacks. As for the second case, various trustworthiness identification and processing schemes have been established to address the issue of low-quality data sources in industrial settings. These schemes encompass sensor node trustworthiness tracing[95], mining association rules using data mining algorithms[96], methods for prognosis of low-quality data integration[97], and data enhancement techniques[98].
III-A2 Data and Network Security Threats in Distributed Industrial Platforms
The distributed architecture of the Industrial Metaverse platform is well-suited for handling large-scale data processing scenarios. However, the introduction of massive P2P links also brings about increased risks of data leakage and management challenges. On one hand, the lifecycle of the Industrial Metaverse spans across different stages, with multiple dispersed stakeholders performing various operations. The increase in digital connections may expose enterprise assets or production processes to data leakage risks. However, leveraging BC-based distributed storage can effectively manage data throughout the PLC. The storage layer provides secure distributed data storage with lightweight and scalability features. Suhail et al.[99] discuss deployment schemes for BC-based Metaverse platforms. They utilized multiple sensors with overlapping perspectives for cross-validation to ensure the credibility of on-chain data. Empowering decision-making through BC in various industrial use cases addresses challenges such as different data repositories and insecure propagation. Peng et al.[100] further consider edge offloading in industrial cloud-edge collaborative architectures, utilizing DT to capture temporal features of computational resources. Additionally, in the second incentive phase, they integrated privacy-security investments to assist industrial agents in making offloading decisions. Shen et al.[101] integrate distributed storage and cloud storage, encrypting raw virtual data in the cloud, storing transaction records in the BC, and maximizing overall benefits for manufacturers and buyers through optimal sampling selection. However, for industrial sensors with higher sampling rates, limitations of real-time sharing need to be further considered. Furthermore, BC is not a one-size-fits-all solution, as issues such as decision paralysis due to information overload, high energy consumption, and latency may arise. Moreover, the nature of the “black box” from BC compels workers to trust the integrity and fairness of processes without understanding the technical basis.
Another issue arising from the distributed architecture of the Industrial Metaverse is the potential for errors and attacks due to multiple untrusted but interdependent participating entities and the integration of non-intersecting data from multiple sources[102]. For instance, man-in-the-middle attacks or altering communication between users and servers may intercept/modify data packets, compromising the confidentiality, integrity, and availability of the system. Cross-site scripting attacks inject external malicious JavaScript code into websites, enabling attackers to inject malicious JavaScript code directly into client websites or have client websites request storage locations. Immersive interaction involves many devices such as wearable devices, headsets, base stations, and controllers, with large-scale data exchange. Despite manufacturers implementing encrypted communication, industrial data may still be accessed by adversaries through eavesdropping methods during transmission. Currently, efforts are underway to enhance industrial network security and anomaly detection. Kim et al.[103] utilize a system based on autoencoders and long short-term memory models to learn normal values of sensors and actuators for monitoring network attacks. Kwon et al.[104] study distributed network protocols widely used in distribution and transmission systems, proposing an anomaly detection method based on bidirectional recurrent networks and IEEE 1815.1. They verified the method with attack data from networked physical systems, successfully detecting eight types of attacks, including unauthorized command attacks and false data injection. However, existing detection systems deployed in Metaverse platforms often result in high computational costs and delays and are difficult to adapt to dynamic attack environments (such as atypical and polymorphic attacks). Therefore, network security researchers must further optimize and promote existing strategies to establish effective models.
III-A3 Inherent Security Vulnerabilities of Enabling Technologies
The enabling technologies mentioned in Section II of the Industrial Metaverse also encompass significant security vulnerabilities that cannot be disregarded.
Concerning AI technology, although Wang et al.[105] highlight its potential for detecting illicit entities and anomalous accounts within the Metaverse, such as discerning suspicious signals through correlating user-generated textual information and spatiotemporal activities, AI also poses various potential risks[106]. On one hand, the industrial platform interconnects a wide array of industrial devices, Metaverse users, collaborative enterprises, and applications, necessitating automated regulation and decision-making due to the intricacies of data management. This grants algorithms extensive operational privileges to reduce managerial costs. However, this approach may give rise to issues like biased decisions[107], low transparency[108], and operational fragility[109]. While certain existing work suggests hierarchical governance structures for algorithm accountability or assesses potential risks of algorithm engines based on AI explainability, a unified compliance governance protocol is yet to be established. On the other hand, due to the black-box nature of AI, malevolent attackers can infiltrate the enterprise source code with malicious code to cause system-level disruption. For instance, they may employ Generative Adversarial Networks (GANs) to fabricate false virtual data or embed malicious software during the model training process, which developers often find challenging to detect in early training results. Therefore, in the industrial context, modular interpretable AI may serve as a promising avenue for future research.
BC records user information and production data in the Industrial Metaverse platform through a distributed ledger and facilitates anonymous P2P interactions through encryption, providing dependable data storage and transmission for enterprises and users[110]. Nevertheless, BC also harbors inherent security risks. Some emanate from the consensus mechanism, such as PoW, wherein an attacker with over half of the BC’s computing power can directly seize control of the entire platform system. Additionally, while consensus-based BC endows robust tamper resistance for industrial data, the computational overhead generated by existing mainstream consensus mechanisms is also a noteworthy concern. Another set of risks arises from BC’s key distribution mechanism. For instance, its decentralized key management implies that private keys might be concentrated and stored on specific entities or servers, rendering all keys vulnerable in the event of an attack on the entity. To avert private key leakage, multi-signature mechanisms can be employed, along with offline storage to mitigate network attacks.
Furthermore, XR-based Industrial Metaverse interaction technologies are susceptible to hacker attacks due to their numerous data interfaces. For example, various systems in factories utilize identity authentication methods such as fingerprint recognition and facial recognition for authorization. Hackers can gain access to device privileges by learning and reconstructing facial information. One immediate countermeasure is the design of novel authentication mechanisms. For instance, Wang et al.[111] devise an identity authentication scheme based on palmprints, which verifies through collecting the reverse scattered signal of acoustic signals, circumventing privacy leakage risks associated with image recognition. Duezguen et al.[112] leverage the concept of semantic connections within the ZeTA protocol to support identity authentication in XR context sharing. Finally, the authors evaluated the correlation between different interaction modes and risks. In the future, adopting compound multi-dimensional authentication technologies, coupled with timely firmware updates, may effectively mitigate privacy breaches during the interaction process. Lastly, beyond the aforementioned issues, addressing security challenges in the Industrial Metaverse necessitates keeping abreast of relevant legal documents and enterprise standards.
III-B Resource Allocation
Due to the increasing number of applications and industrial devices involved in industrial production, distributed deployment of the Metaverse platform may lead to performance imbalances among subsystems and processors, such as computing power and throughput. When facing various types of tasks throughout the PLC, enterprise demands are often dynamic and diverse, posing challenges to the management of network and computing resources. Below, we discuss resource allocation strategies in the Industrial Metaverse focusing on computing resources and storage resources.
III-B1 Computing Resource Allocation Issues
The Industrial Metaverse demands significant computational resources to seamlessly render 3D virtual environments. However, due to the limited computational capabilities of certain lightweight devices within factories, they may fall short for computationally intensive rendering tasks. Although cloud servers offer sufficient computing power for remote rendering, the current state of cloud infrastructure may struggle to meet the ultra-low latency demands of industrial production activities[113].
In collaborative multi-enterprise Industrial Metaverse platforms, some adaptable traditional distributed technologies are gradually being introduced. These include MEC and FL[114]. FL, particularly, is recognized for its privacy-preserving characteristics and excellent scalability as a distributed architecture that aggregates models remotely without exposing data. Al-Quraan et al.[115] discuss the potential applications of FL in various wireless networks, such as visible light communication and massive multiple-input multiple-output (mMIMO), to maintain network performance stability by training models that alleviate microbase station congestion. Nevertheless, despite offloading computation to the edge, FL may still face challenges in frequent information exchange between cloud servers and edge nodes due to the continuous generation of large volumes of new data in industrial production processes. Additionally, the increasing number of parameters and limited bandwidth resources may become bottlenecks for FL. Furthermore, due to differences in device models and performance in factories and varying requirements for model accuracy and computational resources in different scenarios, it is difficult to consider a universal FL aggregation design solution. Furthermore, the number of nodes in Industrial Metaverse systems is often indeterminable during application deployment, necessitating dynamic adjustments in resource allocation to accommodate the addition or removal of nodes[116]. In the face of extreme scenarios, the robustness of the deployment architecture must also be taken into consideration.
Another common architecture is based on the allocation of computing resources in MEC. MEC, by deploying computing nodes near the terminals, is particularly suitable for high-concurrency industrial scenarios. Since edge computing nodes are close to physical devices, they can continue local data processing and decision-making even in cases of unstable or disconnected network connections[117]. Tan et al.[118] introduce DT into edge networks and designed power optimization based on Deep Q-Networks (DQN) for dynamic channel conditions in MEC. The proposed architecture was validated through simulations for applications in automated warehousing logistics and mobile production lines. In Industrial Metaverse platforms, VSPs are responsible for providing resources ordered by various manufacturers. However, excessive resource allocation may lead to resource over-provisioning issues. Ng et al.[119] propose a decision framework for VSPs in the Metaverse based on two-stage stochastic integer programming to minimize overall network costs considering the uncertainty of user demands. However, they did not consider solutions for resource demands in the case of large-scale resource requirements. To address these challenges, Chu et al.[120] decompose the Metaverse platform and utilized application similarity to improve resource utilization. They developed a self-learning framework based on Markov processes to capture implementation characteristics of the Metaverse. Considering the dynamic industrial environment, Yu et al.[121] propose an end-to-end self-organizing resource allocation framework under a 5G heterogeneous network, greatly enhancing Quality of Service (QoS).
In scenarios involving multiple for-profit enterprises, the utility of VSPs in the Industrial Metaverse is considered a function of latency and resource payment costs induced by resource offloading. When the utility is non-negative, VSPs will be willing to invest more computational resources in enterprise collaboration[122]. One scheduling solution involves designing incentive mechanisms to maximize device utilization. Li et al.[123] introduce a distributed crowdsourcing framework called CrowdBC, where smart contracts are used for automated execution, and BC is used for user reputation management and incentive mechanisms. Reputation reflects the past performance of enterprise employees in task resolution. After submitting the execution plan, enterprise employees can request the system to perform task evaluation and record it. Luong et al.[122] propose using multi-agent auctions as an incentive mechanism to support semantic communication (SemCom) and computing resource allocation in the Metaverse. Transmitting semantic information instead of raw data can further reduce offloading latency. Additionally, as the key to synchronizing the real world with the virtual world lies in the data collected by IoT devices and sensors, Han et al.[124] design a dynamic resource allocation framework using a hybrid evolutionary dynamics approach to synchronize the Metaverse with IoT services and data. However, the aforementioned works did not consider the dynamic characteristics of each computing node in the Industrial Metaverse system, such as smart handheld terminals and AGVs. When nodes are predetermined, idle computational resources may remain underutilized when MEC handles computationally intensive tasks.
III-B2 Storage Resource Allocation Issues
Due to the fundamental characteristics of the Metaverse, such as diverse applications, heterogeneous connected devices, ubiquitous user access, and digital identities, in addition to computational capacity, a sustainable storage platform and efficient storage strategy are urgently needed. Servers will collaborate to store information and share content with other servers. In this sense, the payload of redundant data processing and uploading can be reduced by dynamically obtaining terminal information, predicting its demand, and collaboratively storing data, thereby improving the utilization of storage resources and reducing the service latency of the entire system. However, while storing sensor data in the cloud may be a viable solution, as mentioned earlier, cloud-based distributed storage requires continuous internet connectivity and becomes unavailable when the platform encounters emergencies due to internet connection disruptions. Therefore, Rashid et al.[125] proposed the idea of using private edge devices for decentralized distributed data storage in emergencies. They developed a game theory model to dynamically incentivize resource allocation and regulate the QoS of the network. Devices contributing more storage space will receive better QoS (i.e., network bandwidth). To address edge device configuration issues, the authors developed an application that seamlessly runs in Docker containers and connects to the EdgeStore API for system configuration. EdgeStore enables maximizing system availability in emergencies by establishing cooperative storage sharing plans and storing data on private edge devices.
Storage functionality typically includes content storage and policy storage. In the policy storage scheme proposed by Zhao et al.[126], when the server receives relevant requests, it processes tasks based on the request and then returns them to the selected policy device terminal. In Liu et al.’s[127] storage scheme, content is pre-stored on servers. When relevant requests arrive, the server can directly return the data to the terminal. Overall, content storage has lower latency. However, the policy storage scheme can better generalize and adapt to heterogeneous applications. It can support different data requests from users well through preset policies.
Unlike the stable resource scheduling in cloud centers, edge servers are less inclined to share storage space and data due to specific requirements such as workload, latency, and privacy[128]. In emergencies, the inconsistent availability information of storage resources on edge servers may lead to a high data loss rate. To address this issue, Vaquero et al.[129] propose a method to replicate data on all nodes. However, the above method will increase the overall cost of the system and may degrade performance. Recent research work also involves edge caching technology in Metaverse scenarios. Cai et al.[130] design an efficient control policy that integrates computing, caching, and communication (3C) resources and implements the policy by implementing a multi-pipeline traffic control and 3C resource scheduling mechanism. To further optimize the online delivery performance of data-intensive services, the authors also proposed a database placement strategy based on throughput maximization and two effective database replacement strategies. On the other hand, Huynh et al.[131] propose an innovative DT solution that combines MEC and URLLC to support Metaverse applications. From the perspective of the 3C integration model, this solution optimizes edge caching policies, task offloading strategies, and the allocation of computing and communication resources to address the latency optimization problem in the Metaverse enabled by DT.
III-C Interoperability
As a complex system, the Industrial Metaverse platform integrates various sensors and control subsystems with different hardware architectures and control components. If there is no support for interoperability, user identities and industrial equipment data will be limited to specific VSPs, contradicting the original intention of an interconnected Metaverse[132]. Therefore, interoperability of subsystems within the platform and interoperability across platforms are important considerations. Currently, solving the interoperability issues of the Industrial Metaverse mainly starts from two aspects: identity and identification authentication, and interface and integration framework.
Identity and identification define unique codes for users and production equipment, while binding the identification with attributes such as device status and human behavior[133]. This forms the foundation for achieving interoperability. These attributes can be provided by physical devices in factories or created by enterprise personnel themselves. In different platforms, the same identity and identification correspond to the same static attributes. Gadekallu et al.[134] suggest that BC can facilitate the exchange of data between different sub-Metaverses. They proposed using cross-chain protocols to exchange data on multiple BCs located in different virtual worlds. Therefore, Li et al.[133] propose a MetaOpera, a BC-based interoperability protocol for the Metaverse. It supports cross-chain technology and on-chain/off-chain techniques for the interaction between MetaOpera and decentralized platforms as well as top centralized platforms. Compared to authentication technologies based on Sidechains[135], this method achieves lower interaction latency. Additionally, Ghirmai et al.[136] propose using Self-Sovereign Identity (SSI) technology, which replaces traditional identity verification with distributed identity verification using zero-knowledge proofs. For example, two workers docked on a Metaverse platform can prove each other’s public keys according to the principles of SSI and encrypt their interaction records using pre-shared keys. Each worker holds a unique NFT as a unique identifier. However, the authors also pointed out that the issue of BC storage should not be ignored because controlling and managing data and identities may require different storage mechanisms. Additionally, their scenario only applies to static attributes in the Metaverse and does not involve behavioral objects.
Split learning, as a distributed architecture, involves multiple devices and servers in an industrial platform[137]. It can split the backbone network into multiple sub-networks, execute multiple sub-tasks in parallel on different devices or servers, and then integrate the results. Inspired by this, Rawal et al.[138] introduce a Split-Protocol that provides interoperability services. They assume that critical elements and functions are provided by a meta-server, and the system consists of ten identical meta-servers, each executing specific functions of the Metaverse. Role transitions can be achieved through the Split-Protocol. This protocol with role-switching capability can improve the availability of the Metaverse system. Furthermore, Hashash et al.[139] emphasize the coordination between DTs and Edge Metaverse. A Physical Twin (PT) system running in a large-scale sensing area is replicated as a Cyber Twin (CT) system on MEC servers. The area is divided into distributed subgrids and transmitted remotely to MEC servers. Their designed interactive algorithm utilizes optimal transport theory, considering both computational power and data synchronization issues.
In summary, each Industrial Metaverse platform has different configurations, and there is still a lack of standard development patterns. Interoperability and compatibility are guarantees for providing a better operating experience in the Industrial Metaverse. Cooperation between companies and industries is needed to increase the content and value of the Industrial Metaverse and establish an open collaborative ecosystem[140].
TABLE IV: Research on Existing issues in the Industrial Metaverse.
Challenges Issues Ref. Algorithm / Framework Contribution / Performance Year
Security & Confidentiality Sensor-side Industrial Data Source [89] A data misuse prevention mechanism called DAPter. DAPter can substantially raise the bar of the data abuse difficulty with little impact on QoS and overhead. 2021
[91] Kalman anomaly detector based on residuals. This may be helpful for guaranteeing the secure control of a networked system by protecting partial critical sensor measurements from FDI attacks. 2022
[92] A random coding scheme to detect residuals. Estimation-optimization iteration algorithm is to obtain a numerical solution of the coding signal covariance. 2022
Data & Network Security [100] A two-stage incentive mechanism. Efficient resource allocation and computation offloading are realized while privacy is guaranteed. 2023
[101] BC and DT-based distributed storage and cloud storage. The algorithm is better than traditional method for maximizing the total social benefits. 2021
[103] Two payload-based anomaly detection methods. The accuracy and the FP rate are superior to the remaining methods. 2020
[104] An intrusion detection system for an IEEE 1815.1-based power system network. Five types of CMB attacks and three types of FDI and DR attacks are successfully detected. 2020
Inherent Security Vulnerabilities of Enabling Technologies [111] AcoPalm, a palm print recognition system. AcoPalm is resistant to replay and mimicry attacks, and can achieve 96.22% authentication accuracy. 2022
[112] Zeta protocol. The authors explain how it can be used with the available interaction methods offered by head-mounted displays. 2020
Resource Allocation Computing Resource Allocation Issues [118] A framework for MEC based on networked digital dual bases. Web-based DT frameworks with AI capabilities can further reduce task processing latency and improve QoS. 2023
[119] A stochastic optimal resource allocation scheme based on stochastic integer programming. The proposed scheme minimizes the cost of VSPs while taking into account the uncertainty of user demand. 2022
[120] A framework based on semi-Markov decision processes. The proposed approach can realize up to 120% revenue for Metaverse VSPs and a 178.9% probability of acceptance for Metaverse application requests. 2023
[121] Resource allocation mechanism for IIoT in 5G heterogeneous networks. It is possible to achieve better performance than other traditional deep learning (DL) methods and maintain quality of service above accepted levels. 2022
[122] The design of the DL-based auction. Auction improves the revenue while satisfying the individual rationality and incentive compatibility constraints. 2022
[123] A BC-based decentralized crowdsourcing framework. A software prototype is implemented on the Ethernet public test network with real datasets. 2019
[124] A hybrid evolutionary dynamics approach. A dynamic resource allocation framework to synchronize the Metaverse with IoT services and data. 2021
Storage Resource Allocation Issues [125] Edge-based distributed storage system. EdgeStore outperforms state-of-the-art distributed storage systems in terms of throughput, energy consumption, and latency. 2019
[126] Joint optimization of service placement and request routing with multidimensional constraints. The algorithm maximizes the number of requests served by the low latency edge cloud server. 2019
[127] Cache-enhanced multi-user MEC system. The proposed joint optimization of caching, computation, and communication methods can improve energy efficiency at a lower time cost. 2019
[129] Introduced next-generation orchestration technology. Orchestration requirements for next-generation workflow management in microservices architectures and IoT environments are identified. 2018
[130] Designed the first throughput-optimal control strategy. The proposed novel multi-pipeline flow control and 3C resource orchestration mechanism has excellent performance. 2022
Interoperability [141] A protocol called MetaOpera generalized cross-Metaverse interoperability. The size of cross-global proofs and the average time of cross-global transactions are reduced by 8x and 3x, respectively, using the proposed solution compared to the Side-chain solution. 2023
[136] Self-Sovereign Identity (SSI) integrated with BC. Helps address issues of decentralization, trust and interoperability in the Metaverse. 2022
[142] Introduceing an interoperable splitting protocol. The resilience and high-availability of a Metaverse system for role switching. 2022
[139] Iterative algorithms for joint sub-elements and DT associations on MEC servers. The proposed solution coordinates the interactions between the DT and the subuniverse, thus reducing the sub-synchronization time by 25.75%. 2022
IV Standardization
Since the concept of the Metaverse was proposed, efforts have been made to promote its industrial deployment and ensure interoperability and compatibility between different Industrial Metaverse platforms, as well as to reduce the cost of industrial system development and maintenance. Standardization work in this regard has been ongoing.
The MPEG-V standardization, initiated in 2008, defined the exchange of data at the interface between the real world and the virtual world, with the aim of standardizing 4D sensory effects and the virtual world[143]. The standard consists of seven parts, among which 23005.1 provides an overview of the architecture and specifies the relevant information representation for achieving interoperability. 23005.2-3 covers syntax and semantic information, including sensory device functionality and sensor functionality. 23005.5-6 standardize data formats and data types, which are particularly important in industrial multimodal data scenarios[144]. Subsequently, the IEEE Standards Association launched the P2888 project (Interface Networking and Physical World) and established the Synchronization Metaverse as a supplement to the ISO/IEC 23005 series of standards. IEEE 2888 includes four parts: specifications for sensor interfaces in networking and physical worlds, standards for actuator interfaces in networking and physical worlds, orchestration of digital synchronization between networking and physical worlds, and architecture for VR disaster response systems. The first three parts aim to provide general technology for DT or the Metaverse, while the fourth standard aims to provide specific applications for the above three standards[145].
In June 2022, the Metaverse Standards Forum (MSF) was initiated and established by the Khronos Group, with participants including Meta, Microsoft, Epic, NVIDIA, Qualcomm, Sony, and other 37 technology giants engaged in the field of the Metaverse, covering industries such as chip manufacturers and game companies. The organization has now gathered over 2,400 industry vendors and standard organizations to collaborate across the industry on building an open Metaverse, interoperability testing, open-source tools, deployment guidance, and more. In July of the same year, ISO/IEC JTC 1/SC 29 MPEG WG2 Market Need Ad Hoc Group (AHG) launched Metaverse case collection and technical analysis to identify MPEG efficient coding and related system standard technical requirements that can support future Metaverse experiences. In January 2023, the IEC Standardization Evaluation Group 15 on Metaverse (SEG 15) was established. Its specific work includes studying the standardization roadmap for the Metaverse field in conjunction with current standardization achievements and development trends, establishing timely contact with TC/SC/Syc (including JTC1) as well as ISO and other relevant organizations, and providing work proposals for the International Electrotechnical Commission Standards Management Board (IECSMB). Recently, the IEEE P3812 series of standards is in the drafting phase, which defines the requirements for the Metaverse identity framework[146]. The provided identity framework is intended for use across different Metaverse systems to meet the requirements for interoperability/interaction of entities across multiple Metaverse systems. Additionally, the IEEE P2048.101 series of standards define the general technical framework, components, integration, and major business processes for AR systems applied to mobile devices[147].
In the future, as Industrial Metaverse gradually takes off, the accompanying standards regarding interoperability, security and privacy, openness and scalability, cross-industry collaboration, and other aspects will become increasingly refined, forming a comprehensive standard system.
V Future trends and outlook for Industrial Metaverse
While researchers have made significant efforts in studying the Industrial Metaverse tailored to industrial scenarios, alongside notable achievements, there are additional issues warranting consideration. In this section, we further expand our perspective and propose several future research directions for the benefit of researchers. Additionally, for each direction, we provide a brief list of skills/mathematical tools researchers should master, serving as foundational requirements for conducting related research:
V-A Flexible Communication Network Architecture
The seamless internet connectivity of the Metaverse largely relies on reliable broadband services, with holographic communication requiring real-time operation, sufficient bandwidth, and throughput[148]. While low latency and high reliability are the primary advantages of 5G/6G services, consistent latency requires a flexible end-to-end architecture. This is because the Industrial Metaverse requires real-time communication of video, audio, and data. Currently, there is limited research on the network architecture and signal processing aspects of 6G in the context of the Metaverse. Relevant skills/mathematical tools: Wireless communication theory, signal processing, 5G/6G network architecture design.
V-B Privacy and Security Issues
In Section III, we have extensively discussed the confidentiality and security issues of the Industrial Metaverse from the perspectives of data, network, distributed platforms, and enabling technologies, as well as existing solutions. It can be foreseen that, as the Metaverse platform gradually matures, its security measures will largely rely on continuous upgrades through security patches. However, with the increasing prevalence of sophisticated cyber attacks, such security measures may bring challenges in terms of operational costs and scalability[149]. In the future, it is possible to adopt intrinsic security solutions, such as quantum key distribution, which utilizes channel-based keys generated through quantum entanglement to address information leakage in wireless transmissions[102]. Relevant skills/mathematical tools: Quantum key distribution, cryptography, network security.
V-C Lightweight and Ease of Use
Since Industrial Metaverse scenarios are constantly producing, storing, and transmitting data, they require massive access to hardware devices. These devices include production equipment deployed in factories, industrial sensor devices, and access devices such as VR glasses for user interaction. The lightweighting at the hardware level has been slow to materialize. Ways to reduce hardware overhead as well as increase the flexibility and scalability of devices is a valuable topic[150]. Relevant skills/mathematical tools: Embedded system design, hardware optimization, sensor technology.
V-D Model Accuracy and Completeness
The key factor affecting model accuracy in assembly line mode manufacturing applications is the quality of the data collected. The user must provide the system with complete and accurate available data[151]. Although the addition of DT technology can realize virtual samples directly into the model, the data sources often have serious heterogeneity due to different equipment models and data types, and the data fusion and calibration methods at this stage still need to be improved. Relevant skills/mathematical tools: Industrial data preprocessing methods, data fusion methods, machine learning algorithms.
V-E Unified Architecture
Central to the realization of the Metaverse lies the crucial foundation of Web 3.0 architecture. However, the intricate task of amalgamating diverse technologies within a unified architecture necessitates comprehensive efforts. In order to address this challenge, the development of common protocols capable of accommodating the distinct needs and requirements of various industrial sectors becomes imperative. Such protocols would serve as the bedrock for seamless integration, enabling interoperability and synergy among different technologies. By fostering a standardized framework, the development of the Metaverse can progress cohesively, accommodating a multitude of industries and catering to the unique demands of users across diverse domains[152]. Relevant skills/mathematical tools: Network protocol design, distributed system architecture, Web technologies.
V-F Components are Fully Decentralized
The core feature of Web3.0 is decentralization, but APIs at this stage do not integrate well with Web3.0, and new decentralized common protocols need to be developed to achieve complete decentralization of Metaverse components[153]. Only the successful implementation of a decentralized network will occur. This is a challenge because web components and servers still run through legacy controls and protocols, and therefore the backbone infrastructure will need to be extensively changed[154]. Therefore, in the scope of the future, the design of decentralized protocols running on the web needs to be investigated, and open standards and APIs need to be integrated with Web 3.0 engines where data can be seamlessly transferred to different endpoints. Relevant skills/mathematical tools: Decentralized technologies[155], BC development, smart contract programming.
V-G Efficient Cross-chain Authentication
Efficient cross-chain authentication and governance is crucial to ensure the security and legitimacy of digital asset-related activities (e.g., asset transactions) across different submetadomains built on heterogeneous BCs. Current cross-chain mechanisms are mainly focused on digital asset transfers and rely on notary schemes, hash locks, relay chains and side chains, with few mechanisms considering cross-chain authentication and governance for meta-domains. The implementation, efficiency and security of cross-domain and BC authentication need further research. Moreover, novel decentralized, hierarchical and penetrating cross-chain governance mechanisms also need further research in the Metaverse. In addition, efficient meta-space-specific consensus mechanisms, redesigned block structures, and well-designed user incentives are all necessary for unique meta-space applications[156]. In summary, open challenges include application-specific governance rule design, programmable and scalable cross-chain governance architecture design, on-chain entity identification and risk assessment, dynamic and collaborative cross-chain governance, etc. Relevant skills/mathematical tools: BC cross-chain technology, authentication protocols, smart contract programming.
V-H Integration with Large-scale Pre-trained Models
Large-scale pre-trained models hold the potential to offer exquisitely tailored and individualized encounters within the Metaverse. Through discerning analysis of user inclinations, behaviors, and historical data, these models can be refined to generate bespoke virtual panoramas[157], characters, and interactions, ushering in a realm of heightened immersion and personalized engagement within the Metaverse.
Nevertheless, notable challenges persist. One such challenge lies within the realm of the language macrocosm, where Metaverse users hail from diverse nations and regions, conversing in various languages. Consequently, it becomes imperative for the model to possess the capability to seamlessly interact across linguistic boundaries. Furthermore, the endeavor to ensure the seamless deployment of the model, endowed with high concurrency and unwavering stability, becomes an ineluctable quandary[158]. This ensures swift and responsive handling of users’ requests, a matter of utmost significance. Relevant skills/mathematical tools: Natural language processing, deep learning model training, model deployment and optimization.
V-I Integration with Quantum Computing
Quantum computing, with its unparalleled capacity to handle intricate calculations and tackle problems exponentially faster, possesses the potential to reshape the capabilities of the Metaverse and elevate its performance to unprecedented heights. A recent study conducted by Cui et al.[159] delves into the implementation of secure communication and efficient cross-chain protocols utilizing quantum computing. Meanwhile, Ren et al.[160] have explored the utilization of game-based quantum collective learning and many-to-many matching schemes within the Metaverse, enabling the attainment of optimal strategies for maximizing system revenue. Looking ahead, quantum computing is poised to find widespread application in enhancing data processing, encryption, and security performance within the Metaverse. Furthermore, it holds the promise of expediting virtual and AR experiences, thereby propelling the Metaverse into a realm of heightened immersion and seamless interactivity. Relevant skills/mathematical tools: Quantum computing theory, quantum algorithm design, quantum communication protocols.
V-J Availability of Datasets for Industrial Metaverse
In industrial environments, there is a plethora of data in different types, formats, and sources. These data often reside in disparate systems and devices, leading to data isolation and fragmentation. Lack of unified data integration and management mechanisms hampers seamless integration and consolidation of data. Moreover, various stakeholders in industrial environments may possess different data ownership and access rights. Establishing appropriate mechanisms for data ownership and access control is essential to ensure data legitimacy and compliance. Relevant skills/mathematical tools: Data management, data integration, access control.
V-K Ethical and Social Impacts
When machine decisions impact people’s lives, especially in industrial scenarios with high cost and safety requirements, it is important to clarify how much autonomy the system allocates to machine algorithms. In addition to the technical aspects, fairness in resource allocation must also be considered[161]. For example, some large companies may find it easier to concentrate high-quality and extensive digital resources for DT modeling, while expensive hardware may exclude organizations with lower purchasing power, exacerbating the digital divide. It is necessary to enact more artificial intelligence laws in the future to subject the operation of the industrial Metaverse to legal jurisdiction.
VI Conclusion
In this paper, we conducted comprehensive research on the Industrial Metaverse. By analyzing the characteristics of Metaverse technology, we put forward convincing reasons for deploying Metaverse platforms in the industrial field. Furthermore, we summarized the research progress of several key enabling technologies in the Industrial Metaverse and elucidated the advantages of each technology in various stages of industrial production. We provided a comprehensive summary of various technologies, including existing challenges and performance requirements in industrial scenarios. Furthermore, starting from the Metaverse itself and combining with the characteristics of industrial scenarios, we identified three key challenges: security and confidentiality issues, resource allocation problems under resource constraints, and interoperability. In addition, we compiled the standardization efforts in the field of Metaverse over the past few years. Lastly, we concluded with a forward-looking discussion, considering the current challenges and future research directions. With the continuous advancement of Metaverse-related technologies and their further integration with industrial scenarios, we anticipate that our work can provide references for future research endeavors.